Solve

ut = uxx 0<x<l, t>0

u(0, t)= u(l, t)= 0

u(x, 0)= x(l-x) , 0<=x<=l



1
Expert's answer
2021-02-04T08:16:04-0500

Solution

Solution can be found in the form  u(x,t) = X(x)*T(t)

X’’/X = T’/T  => X’’/X = T’/T = -λ2  

X’’+ λ2X = 0  =>  X = Asin(λ*x) + Bcos(λ*x) . Here A,B are arbitrary constants.

From boundary condition u(0, t) = 0  => B = 0  

From boundary condition u(l, t) = 0  => sin(λ*x) = 0  =>  λ = π*n/l (n = 1,2, …)  

So  X = Asin(π*n *x/l)

Equation on T gives now  T’ + λ2 T = 0  =>  T = exp(- λ2 t) = exp(- (π*n/l)2 t)   

And un = An* exp(- (π*n/l)2 t)* sin(π*n *x/l)

"u(x,t)=\displaystyle\sum_{n=1}^\infty A_n e^{-(\pi n/l)^2t} sin(\pi n \frac{x}{l})"

From initial condition u(x, 0)= x(l-x) , 0<=x<=l

"x(l-x)=\displaystyle\sum_{n=1}^\infty A_n sin(\pi n \frac{x}{l})"

"A_n=\frac{\displaystyle\int_0^l x(l-x) sin(\pi n x/l) dx}{\displaystyle\int_0^l sin^2(\pi n x/l) dx}"

"A_n=
\frac{2}{l} \displaystyle\int_0^l x(l-x) sin(\pi n x/l) dx"


Let y = x/l => "A_n=
2l^2 \displaystyle\int_0^1 y(1-y) sin(\pi n y) dy"

Let z = π*n *y  => "A_n=
\frac{2l^2}{(\pi n)^2} \displaystyle\int_0^{\pi n} z(1-\frac{z}{\pi n}) sin(z) dz"


"A_n=\frac{2l^2}{(\pi n)^2} [sin(z)-z cos(z)
-\frac{1}{\pi n} (2z sin(z)-(z^2-2) cos(z))] |_0^{\pi n}"

"A_n\u200b=\\frac{2l^2}{(\\pi n)^2} [\\pi n (-)^{n+1}+\\frac{1}{\\pi n}((\\pi n)^2-2)(-)^n+\\frac{2}{\\pi n}]"

"A_n\u200b=\\frac{4l^2}{(\\pi n)^3} [(-)^{n+1}+1]"

A2m = 0, A2m+1 = 8l2/(π*(2m+1))3   (m = 0,1, …)

"u(x,t)=8l^2\displaystyle\sum_{m=0}^\infty \frac{1}{(\pi (2m+1))^3} e^{-(\pi (2m+1)/l)^2t} sin(\pi (2m+1) \frac{x}{l})"


Answer


"u(x,t)=8l^2\displaystyle\sum_{m=0}^\infty \frac{1}{(\pi (2m+1))^3} e^{-(\pi (2m+1)/l)^2t} sin(\pi (2m+1) \frac{x}{l})"